Solve equations by factoring
Solve the equation.
(x - 8)^2 = - 36x

Answer by ewatrrr(24785)


Hi
Solving equation by factoring
(x - 8)^2 = - 36x
x^2 - 16x + 64 = -36x
x^2 + 20x + 64 = 0
(x + 4)(x+ 16) = 0 Note:SUM of the inner product(4x) and the outer product(16x) = 20x
(x + 4)=0
x = -4
(x+ 16) = 0
x = -16